### v1

Abstract
Divergence constraints are present in the governing equations of many physical phenomena, and they usually lead to a Poisson equation whose solution typically is the main bottleneck of many simulation codes. Algebraic Multigrid (AMG) is arguably the most powerful preconditioner for Poisson’s equation, and its effectiveness results from the complementary roles played by the smoother, responsible for damping high-frequency error components, and the coarse-grid correction, which in turn reduces low-frequency modes. This work presents several strategies to make AMG more compute-intensive by leveraging reflection, translational and rotational symmetries, often present in academic and industrial configurations. The best-performing method, AMGR, is based on a multigrid reduction framework that introduces an aggressive coarsening to the multigrid hierarchy, reducing the memory footprint, setup and application costs of the top-level smoother. While preserving AMG’s excellent convergence, AMGR allows replacing the standard sparse matrix-vector product with the more compute-intensive sparse matrix-matrix product, yielding significant accelerations. Numerical experiments on industrial CFD applications demonstrated up to 70% speed-ups when solving Poisson’s equation with AMGR instead of AMG. Additionally, strong and weak scalability analyses revealed no significant degradation.
keywords:
AMG, Multigrid reduction, SpMM, Spatial symmetries, Poisson’s equation
{MSCcodes}
65F08, 65F50, 65N55, 65Y05
1 Introduction
Divergence constraints are prevalent in physical problems, often following fundamental conservation principles like mass or electrical charge conservation. Such constraints lead to a Poisson equation that plays a fundamental role in many areas of science and engineering, such as computational fluid dynamics (CFD), linear elasticity, and electrostatics. Indeed, the solution of the associated linear system is generally the most computationally intensive part of scientific simulation codes, and the design and implementation of Poisson solvers are far from straightforward due to the interplay of numerical and hardware challenges.
Historically, the most efficient way to solve Poisson’s problems is through iterative methods based on Krylov subspaces [7, 46, 54], whose implementation is simple and easily parallelisable, requiring only basic linear algebra operations. Namely, matrix by vector products, scalar products and vector updates. However, iterative linear solvers must be properly preconditioned to be effective [47, 10]. The choice, design and implementation of such preconditioners are not trivial, and it is one of the most active research fields in numerical analysis. Preconditioners based on incomplete factorisations were very popular in the early days of numerical linear algebra. The first papers suggesting this approach were [39, 19], where the factorisation takes place with no fill-in, i.e., the pattern of the Cholesky factor, $L$, equals the lower pattern of $A$, with $A\simeq LL^{T}$. Subsequently, more advanced and effective alternatives with dynamic fill-in control were proposed; see, for instance, [45, 37].
However, the sequential nature of such methods and the increasing availability of parallel computers in the late ’90s made them lose ground against other alternatives with higher degrees of parallelism. For instance, the application of preconditioners based on approximate inverses solely relies on the sparse matrix-vector product (SpMV), an easily parallelisable operation. In some cases, even their construction is reasonably concurrent. The most prominent variants are AINV [12, 13, 11], SPAI [25] and the Factored Sparse Approximate Inverse (FSAI) [33, 31]. While approximate inverses provide a high degree of parallelism, they are not optimal in the sense that when the mesh size decreases (hence increasing the linear system size), the problem becomes more ill-conditioned and more iterations are required to reach the same accuracy.
This problem worsens nowadays, as extreme-scale linear systems must be solved on massively parallel supercomputers, and single-level preconditioners such as incomplete factorisations or approximate inverses generally require excessive iterations. The problem of scalability is overcome with multilevel preconditioners like Geometric Multigrid (GMG) or Algebraic Multigrid (AMG) ][52, 55, 59]. Thanks to the interplay between smoother and coarse-grid correction, these methods often solve a given PDE with a number of iterations independent of the mesh size. Many freely available multigrid packages exist. To cite a few, Hypre [28], Trilinos [14], and PETSc [9] provide very effective implementations with excellent scalability.
To develop efficient and scalable solvers, it is necessary to identify the limitations of current computing devices and develop algorithms that overcome them. For instance, the low arithmetic intensity of most sparse linear algebra kernels motivated strategies like using mixed precision [8] or applying more compute-intensive algorithms [29]. Similarly, the large memory to network bandwidth ratio led to implementations that hide or completely avoid inter-node communications [34, 42, 48]. Additionally, the limited available memory resulted in approaches like exploiting data sparsity [23, 6].
A few works exploiting symmetries for solving Poisson’s equation exist [5, 4, 24, 49]. This paper extends them by proposing enhanced variants of FSAI and AMG preconditioners. Their advantages result from mitigating several of the computational challenges above. Namely, we show that given an arbitrarily complex geometry presenting reflection, translational or rotational symmetries, it is possible to apply a consistent ordering that makes the coefficient matrix (and preconditioners) satisfy regular block structures. This, in turn, allows the standard SpMV to be replaced with the more compute-intensive sparse matrix-matrix product (SpMM[3], leading to lighter and faster variants of FSAI and AMG. Several strategies are considered, the best performing one based on a multigrid reduction framework that induces an aggressive coarsening to the multigrid hierarchy, reducing the memory footprint, setup and application costs of the top-level smoother. While preserving the excellent convergence of the standard AMG, replacing SpMV with SpMM yielded significant speed-ups.
Without loss of generality, the targeted applications are incompressible CFD simulations. In particular, direct numerical simulation (DNS) and large-eddy simulation (LES) of turbulent flows, which are essential for many areas of engineering, such as energy production and environmental monitoring. In such areas, spatial symmetries are commonly exhibited. For instance, most vehicles have a central reflection symmetry [1, 16]. Cases with two reflection symmetries are also usual, including jets [20] and building simulations [41]. Further gains can be attained by exploiting translational or rotational symmetries. They are present in turbomachinery [57], heat exchangers [38, 44], nuclear reactors [21, 40], or recurring structures, such as arrays of buildings [26] or wind farms [15].
The remaining sections are organised as follows. Section 2 presents a strategy for exploiting reflection symmetries through low-rank corrections, and Section 3 extends it to translational and rotational symmetries. Section 4 discusses the parallel implementation, Section 5 presents meaningful numerical experiments, and Section 6 gives some concluding remarks.
2 Symmetry-aware spatial discretisation
Let us start by recalling a strategy to enhance FSAI and later extend it towards AMG. Such a strategy introduces low-rank corrections to enable the use of SpMM but requires a discretisation consistent with the spatial symmetries. Given an arbitrary mesh presenting a single reflection symmetry, we will order its grid points by first indexing the ones lying on one half and then those on the other. Therefore, analogous to Fig. 1, if we impose the same local ordering (mirrored by the symmetry’s hyperplane) to the resulting two subdomains, we ensure that all the scalar fields satisfy:
(1) $x=\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}\in\mathbb{R}^{n},$
where $n$ stands for the mesh size and $x_{1},x_{2}\in\mathbb{R}^{n/2}$ for $x$’s restriction to each of the subdomains. Mirrored grid points are in the same position within the subvectors, and discrete versions of virtually all partial differential operators satisfy the following block structure:
(2) $H=\begin{pmatrix}H_{1,1}&H_{1,2}\\ H_{2,1}&H_{2,2}\end{pmatrix}\in\mathbb{R}^{n\times m},$
where $H_{i,j}\in\mathbb{R}^{n/2\times m/2}$ accounts for the couplings between the $i$th and $j$th subdomains. As long as $H$ only depends on geometric quantities (which is typically the case) or on material properties respecting the symmetries, given that both subdomains are identical and thanks to the mirrored ordering, we have that:
(3) $H_{1,1}=H_{2,2}\text{ and }H_{1,2}=H_{2,1},$
and, by denoting $H_{i}\equiv H_{1,i}$, we can rewrite Eq. 2 as:
(4) $H=\begin{pmatrix}H_{1}&H_{2}\\ H_{2}&H_{1}\end{pmatrix}.$
Figure 1: Single-symmetry 1D mesh with a random mirrored ordering.
The procedure above can be applied recursively to exploit an arbitrary number of reflection symmetries, $s$. For instance, taking advantage of $s=2$ symmetries results in 4 mirrored subdomains and, analogously to Eq. 4, virtually all discrete operators satisfy the following:
(5) $H=\begin{pmatrix}H_{1}&H_{2}&H_{3}&H_{4}\\ H_{2}&H_{1}&H_{4}&H_{3}\\ H_{3}&H_{4}&H_{1}&H_{2}\\ H_{4}&H_{3}&H_{2}&H_{1}\end{pmatrix},$
where $H_{i}\in\mathbb{R}^{n/4\times m/4}$ contains the couplings between the first and $i$th subdomains.
Thanks to the discretisation presented above, exploiting $s$ reflection symmetries allows meshing a $1/2^{s}$ fraction of the entire domain, henceforth named base mesh. Figure 3 depicts a schematic representation of the base mesh on an arbitrary domain with two symmetries. Then, instead of building the entire operators, $H\in\mathbb{R}^{n\times m}$, it is only needed to build the base mesh’s couplings with itself, $H_{1}\in\mathbb{R}^{n/2^{s}\times m/2^{s}}$, and with its $12^{s}-$ 1 mirrored counterparts, $H_{2},\dots,H_{2^{s}}\in\mathbb{R}^{n/2^{s}\times m/2^{s}}$. As a result, both the setup and memory footprint of the matrices are reduced by a factor of $2^{s}$ [3].
Furthermore, while the sparsity pattern of $H_{1}$ matches that of the actual operator built upon the base mesh, the outer-subdomain couplings, $H_{2},\dots,H_{2^{s}}$, have very few non-zero entries (if any), making the following splitting very advantageous:
(6) $H=\mathbb{I}_{2^{s}}\otimes H_{\text{inn}}+H_{\text{out}},$
where $H_{\text{inn}}\coloneqq H_{1}\in\mathbb{R}^{n/2^{s}\times m/2^{s}}$ and $H_{\text{out}}\coloneqq H-\mathbb{I}_{2^{s}}\otimes H_{1}\in\mathbb{R}^{n\times m}$.
The splitting of Eq. 6 allows the standard SpMV by $H$ to be replaced with a specialised version of the more compute-intensive SpMM. The resulting kernel is a fusion of an SpMM by $H_{\text{inn}}$, an SpMV by $H_{\text{out}}$ and a linear combination of vectors (axpy), and has been employed throughout the simulations to make up to 5x faster their matrix multiplications [3]. In order to elucidate the computational advantages of SpMM, let us note that the standard approach for applying $\mathbb{I}_{2^{s}}\otimes H_{\text{inn}}$ to a vector is through an SpMV call, which performs the following operation:
(7) $y=\begin{pmatrix}H_{\text{inn}}&&\\ &\ddots&\\ &&H_{\text{inn}}\end{pmatrix}\begin{pmatrix}x_{1}\\ \vdots\\ x_{2^{s}}\end{pmatrix}\in\mathbb{R}^{n}.$
Then, replacing SpMV with SpMM algebraically corresponds to:
(8) $(y_{1}\dots y_{2^{s}})=H_{\text{inn}}(x_{1}\dots x_{2^{s}})\in\mathbb{R}^{n/2^{s}\times 2^{s}}.$
The fact that SpMM reads $H_{\text{inn}}$ $2^{s}$ fewer times makes its arithmetic intensity considerably higher and, since SpMV and SpMM are generally memory-bound kernels, this increase translates into significant speed-ups.
Apart from accelerating matrix multiplications, symmetries can be further harnessed to decompose Poisson’s equation into a set of decoupled subsystems. For the sake of clarity, let us recall the single-symmetry case of Fig. 1. Following Eq. 6, the discrete Laplacian can be written as:
(9) $A=\begin{pmatrix}A_{1}&A_{2}\\ A_{2}&A_{1}\end{pmatrix}\in\mathbb{R}^{n\times n},$
and, by denoting the identity matrix with $\mathbb{I}_{k}\in\mathbb{R}^{k\times k}$, we can define the following change-of-basis:
(10) $Q_{1}\coloneqq\frac{1}{\sqrt{2}}\left(\begin{array}[]{c c}\mathbb{I}_{\nicefrac{{n}}{{2}}}&\mathbb{I}_{\nicefrac{{n}}{{2}}}\\ \mathbb{I}_{\nicefrac{{n}}{{2}}}&-\mathbb{I}_{\nicefrac{{n}}{{2}}}\end{array}\right)\in\mathbb{R}^{n\times n},$
which transforms $A$ into:
(11) $Q_{1}AQ_{1}^{-1}=\left(\begin{array}[]{c c}A_{1}+A_{2}&\\ &A_{1}-A_{2}\end{array}\right).$
As shown in [5], the block diagonalisation above can be generalised to an arbitrary number of reflection symmetries, $s$, by defining the following change-of-basis:
(12) $Q_{s}\coloneqq\prod_{i=1}^{s}\left(\mathbb{I}_{2^{i-1}}\otimes\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}\otimes\mathbb{I}_{\nicefrac{{n}}{{2^{i}}}}\right)\in\mathbb{R}^{n\times n},$
which satisfies $Q_{s}^{-1}=Q_{s}$ and transforms the discrete Laplacian into $2^{s}$ subsystems:
(13) $\hat{A}\coloneqq Q_{s}AQ_{s}^{-1}=\begin{pmatrix}\hat{A}_{1}&&\\ &\ddots&\\ &&\hat{A}_{2^{s}}\end{pmatrix}.$
Then, similarly to Eq. 11, $\hat{A}$ can be split as follows:
(14) $\hat{A}=\mathbb{I}_{2^{s}}\otimes A_{\text{inn}}+\begin{pmatrix}A_{\text{out},1}&&\\ &\ddots&\\ &&A_{\text{out},2^{s}}\end{pmatrix},$
making the Poisson solver of Algorithm 1 compatible with SpMM. Additionally, the decoupled solution of $\hat{A}$’s subsystems in line 4 makes Krylov subspace methods converge faster [5].
Algorithm 1 Poisson solver exploiting $s$ reflection symmetries
1:$\hat{A}_{1},\dots,\hat{A}_{2^{s}}$, $Q_{s}$ and $b\in\operatorname{range}(A)\subseteq\mathbb{R}^{n}$
2:procedure Solve($b$)
3:     Transform forward: $\hat{b}=Q_{s}b$
4:     Decoupled solution of $\hat{A}_{i}\hat{x}_{i}=\hat{b}_{i}\;\;\forall i\in\{1,\dots,2^{s}\}$
5:     Transform backward: $x=Q_{s}\hat{x}$
6:     return $x$
7:end procedure
2.1 Low-rank corrections for factorable preconditioners
In this section, we aim to develop low-rank corrections for enhancing factorable preconditioners, specifically focusing on making FSAI compatible with SpMM. Although initially introduced in [2], our current objective is to extend this strategy towards non-factorable preconditioners like AMG.
The idea of applying low-rank corrections arises from the close similarity between $\hat{A}$’s subsystems. Indeed, in Eq. 14, all the outer-couplings are substantially sparser than the inner and, as discussed in [2], $\operatorname{rank}(A_{\text{out},i})=\mathcal{O}(n^{2/3})$, whereas $\operatorname{rank}(A_{\text{inn}})=\mathcal{O}(n)$. Hence, for all $i\in\{1,\dots,2^{s}\}$:
(15) $\operatorname{rank}\left(A_{\text{out},i}\right)\ll\operatorname{rank}\left(A_{\text{inn}}\right),$
and it makes sense to introduce another level of approximation to FSAI by assuming that each of $\hat{A}$’s subsystems satisfies:
(16) $\hat{A}_{i}=A_{\text{inn}}+A_{\text{out},i}\simeq A_{\text{inn}}.$
In the context of preconditioning linear systems, much work has recently been devoted to low-rank matrix representations [35, 23, 36, 22, 58]. Let us recall the following result [36].
Theorem 2.1.
Given the two SPD matrices $A$ and $B$, let $L$ be the lower Cholesky factor of $B$, i.e., $B=LL^{T}$. Then, given $Y\coloneqq(\mathbb{I}-L^{-1}AL^{-T})$, the following holds:
$A^{-1}=B^{-1}+L^{-T}V\Sigma\left(\mathbb{I}-\Sigma\right)^{-1}V^{T}L^{-1},$
where $Y=Y^{T}$ and $Y=V\Sigma V^{T}$ is the eigendecomposition of $Y$.
Proof 2.2.
From the definition of $Y$ it follows that $(\mathbb{I}-Y)^{-1}=L^{T}A^{-1}L$ and, after some straightforward calculations:
$A^{-1}=B^{-1}+L^{-T}Y(\mathbb{I}-Y)^{-1}L^{-1}.$
Let $V\Sigma V^{T}$ be the eigendecomposition of $Y$. Then:
$(\mathbb{I}-Y)^{-1}=V(\mathbb{I}-\Sigma)^{-1}V^{T}$
and, therefore, $A^{-1}=B^{-1}+L^{-T}V\Sigma\left(\mathbb{I}-\Sigma\right)^{-1}V^{T}L^{-1}$.
At this point, let us consider the FSAI of $A_{\text{inn}}$. It provides an approximation to the inverse of $A_{\text{inn}}$’s lower Cholesky factor, $G_{\text{inn}}\simeq L_{\text{inn}}^{-1}$, ensuring that:
(17) $G_{\text{inn}}^{T}G_{\text{inn}}\simeq A_{\text{inn}}^{-1}.$
Then, for each subsystem $\hat{A}_{i}$, we can define the following auxiliary matrix:
(18) $Y\coloneqq\mathbb{I}_{\nicefrac{{n}}{{2^{s}}}}-G_{\text{inn}}\hat{A}_{i}G_{\text{inn}}^{T}\in\mathbb{R}^{n/2^{s}\times n/2^{s}},$
and, by virtue of Theorem 2.1, we have that:
(19) $\hat{A}_{i}^{-1}=G_{\text{inn}}^{T}G_{\text{inn}}+G_{\text{inn}}^{T}V\Sigma\left(\mathbb{I}_{\nicefrac{{n}}{{2^{s}}}}-\Sigma\right)^{-1}V^{T}G_{\text{inn}}.$
The fact that $Y$’s eigendecomposition is dense prevents applying the “full-rank” correction of Eq. 19. However, thanks to Eq. 15, we can expect $Y$ to have a high data sparsity, i.e.,  its action is well represented by a low-rank approximation that only accounts for its $k$ most relevant eigenpairs:
(20) $Y\simeq V_{k}\Sigma_{k}V_{k}^{T},$
where $V_{k}\in\mathbb{R}^{n/2^{s}\times k}$ and $\Sigma_{k}\in\mathbb{R}^{k\times k}$ yield the following low-rank correction:
(21) $\hat{A}_{i}^{-1}\simeq G_{\text{inn}}^{T}G_{\text{inn}}+Z_{k}\Theta_{k}Z_{k}^{T},$
with $Z_{k}\coloneqq G_{\text{inn}}^{T}V_{k}\in\mathbb{R}^{n/2^{s}\times k}$ and $\Theta_{k}\coloneqq\Sigma_{k}(\mathbb{I}_{k}-\Sigma_{k})^{-1}\in\mathbb{R}^{k\times k}$.
As for selecting the most relevant eigenpairs, it is enough to remark that $Y$ measures how far each preconditioned subsystem, $G_{\text{inn}}\hat{A}_{i}G_{\text{inn}}^{T}$, is from the identity matrix. Then, given the harmful effect that small eigenvalues have in the Preconditioned Conjugate Gradient (PCG) convergence [53], $Y$’s most effective eigenpairs are those associated with the smallest eigenvalues of $X\coloneqq G_{\text{inn}}\hat{A}_{i}G_{\text{inn}}^{T}$. Hence, by computing a low-rank approximation of $X$:
(22) $X\simeq U_{k}\Lambda_{k}U_{k}^{T},$
we can obtain the following truncated eigendecomposition of $Y$:
(23) $Y\simeq U_{k}(\mathbb{I}_{k}-\Lambda_{k})U_{k}^{T}.$
It can be shown experimentally that rough and cost-effective approximations of $U_{k}$ and $\Lambda_{k}$ suffice for preconditioning purposes. Then, applying the above procedure to each of the $2^{s}$ subsystems leads to the following low-rank corrected FSAI, henceforth denoted as LRCFSAI($k$):
(24) $\mathbb{I}_{2^{s}}\otimes G_{\text{inn}}^{T}G_{\text{inn}}+\begin{pmatrix}Z_{k,1}\Theta_{k,1}Z_{k,1}^{T}&&\\ &\ddots&\\ &&Z_{k,2^{s}}\Theta_{k,2^{s}}Z_{k,2^{s}}^{T}\end{pmatrix}.$
LRCFSAI($k$) is compatible with SpMM and, thanks to reusing the same FSAI on all the subsystems, grants savings in the preconditioner’s memory requirements and setup costs by a factor of $2^{s}$. Of course, this comes at the price of using lower quality approximations, given that $G_{\text{inn}}$ does not account for the outer-couplings. However, introducing low-rank corrections proved very effective despite representing a relatively low overhead [2]. It is remarkable that applying a rank-$k$ correction to each of the $2^{s}$ subsystems separately corresponds to applying a rank-$(2^{s}k)$ correction on the global system, $\hat{A}$.
In order to illustrate the behaviour of LRCFSAI($k$) and, most especially, to assess the quality of the AMG variants that we will develop in the following sections, let us consider the following Poisson’s equation with homogeneous Neumann boundary conditions:
(25) $-\frac{\partial^{2}u}{\partial x^{2}}-\frac{\partial^{2}u}{\partial y^{2}}-\frac{\partial^{2}u}{\partial z^{2}}$ $=f\text{ in }\Omega$
$\frac{\partial u}{\partial n}$ $=0\text{ on }\partial\Omega$
[…]
(26) $x_{i}=\frac{1}{2}\left(1+\frac{\tanh\left(\gamma_{x}\left(2\frac{(i-1)}{n_{x}}-1\right)\right)}{\tanh\left(\gamma_{x}\right)}\right)\;\;\forall i\in\{1,\dots,n_{x}+1\},$
analogously applied in the $y$- and $z$-directions using $\gamma_{x}=\gamma_{y}=\gamma_{z}=1.5$.
Table 1 summarises the convergence of LRCFSAI($k$) on the model problem of Eq. 25. Clearly, low-rank corrections outweigh the fact of ignoring the outer-couplings, actually making LRCFSAI($k$) converge faster than the standard FSAI. As a result, we proceeded with its parallel implementation, which is out of the scope of this work but was thoroughly reviewed in [2].
Table 1: PCG + LRCFSAI($k$) results on the model problem with $n=64^{3}$.
iterations
preconditioner $s=0$ $s=1$ $s=2$ $s=3$
FSAI 291 220 173 121
LRCFSAI(0) 291 241 193 149
LRCFSAI(1) 228 194 146 106
LRCFSAI(2) 228 172 124 92
LRCFSAI(4) 186 136 107 83
LRCFSAI(8) 131 111 81 61
LRCFSAI(16) 113 82 63 50
2.2 Low-rank corrections for non-factorable preconditioners
At this point, we can extend the strategy presented in Section 2.1 towards AMG, which is probably the most powerful preconditioner for Poisson’s equation. The effectiveness of AMG is given by the complementary roles played by the smoother, which is responsible for damping high-frequency error components, and the coarse-grid correction, which in turn reduces low-frequency modes. Large problems require a progressive coarsening into a hierarchy of smaller and smaller grids. These grids are created using the concept of strength of connection, which measures the likelihood that the smooth error components on two adjacent nodes have similar values. There are a number of beautiful books on the subject, e.g.[51, 55].
Bearing in mind that the application of AMG relies on matrix multiplications, we aim to accelerate it by replacing SpMV with SpMM. However, AMG is not explicitly factorable, and we need to develop alternative low-rank corrections. With this aim, let us recall the following result [36].
Theorem 2.3.
Given the two SPD matrices $A$ and $B$, let us define the auxiliary matrix $Y\coloneqq(\mathbb{I}-B^{-1}A)$. Then, the following holds:
$A^{-1}=\left(\mathbb{I}+U\left(\Sigma^{-1}-V^{T}U\right)^{-1}V^{T}\right)B^{-1},$
where $Y\neq Y^{T}$ and $Y=U\Sigma V^{T}$ is the eigendecomposition of $Y$.
Proof 2.4.
From the definition of $Y$, it follows that $A^{-1}=(\mathbb{I}-Y)^{-1}B^{-1}$. Then, applying the Sherman-Morrison-Woodbury formula, we have that:
$(\mathbb{I}-Y)^{-1}=(\mathbb{I}-U\Sigma V^{T})^{-1}=\mathbb{I}+U(\Sigma^{-1}-V^{T}U)^{-1}V^{T},$
and combining both equations completes the proof.
As we did to derive LRCFSAI($k$), let us start by considering the following AMG approximation of $A_{\text{inn}}^{-1}$:
(27) $M_{\text{inn}}\simeq A_{\text{inn}}^{-1}.$
Then, for each subsystem $\hat{A}_{i}$, we can define the following auxiliary matrix:
(28) $Y\coloneqq\mathbb{I}_{\nicefrac{{n}}{{2^{s}}}}-M_{\text{inn}}\hat{A}_{i}\in\mathbb{R}^{n/2^{s}\times n/2^{s}},$
and, by virtue of Theorem 2.3, we have that:
(29) $\hat{A}_{i}^{-1}=\left(\mathbb{I}_{\nicefrac{{n}}{{2^{s}}}}+U\left(\Sigma^{-1}-V^{T}U\right)^{-1}V^{T}\right)M_{\text{inn}}.$
Once again, thanks to Eq. 15, we can expect $Y$ to be well represented by a truncated eigendecomposition only accounting for its $k$ most relevant eigenvectors:
(30) $Y\simeq U_{k}\Sigma_{k}V_{k}^{T},$
where $U_{k},V_{k}\in\mathbb{R}^{n/2^{s}\times k}$ and $\Sigma_{k}\in\mathbb{R}^{k\times k}$. Differently to Eq. 20 in Section 2.1, $Y$’s nonsymmetry makes it require both the right- and left-eigenvectors, yielding the following low-rank correction:
(31) $\hat{A}_{i}^{-1}\simeq\left(\mathbb{I}_{\nicefrac{{n}}{{2^{s}}}}+U_{k}(\Sigma_{k}^{-1}-V_{k}^{T}U_{k})^{-1}V_{k}^{T}\right)M_{\text{inn}}=U_{k}\Theta_{k}V_{k}^{T}M_{\text{inn}},$
where we considered biorthonormal bases for the right- and left-eigenvectors, i.e., $V_{k}^{T}U_{k}=\mathbb{I}_{k}$, and defined $\Theta_{k}\coloneqq\mathbb{I}_{k}+(\Sigma_{k}^{-1}-\mathbb{I}_{k})^{-1}\in\mathbb{R}^{k\times k}$. Then, by recalling the Sherman-Morrison-Woodbury formula, we have that: $\Theta_{k}=\mathbb{I}_{k}+\Sigma_{k}+\Sigma_{k}(\mathbb{I}_{k}-\Sigma_{k})^{-1}\Sigma_{k}$.
When it comes to selecting the most relevant eigenvectors, let us note that, as in Section 2.1, $Y$ measures how far it is each preconditioned subsystem, $M_{\text{inn}}\hat{A}_{i}$, from the identity matrix. Then, given the harmful effect of small eigenvalues in the convergence of Krylov subspace methods, $Y$’s most effective eigenvectors are those associated with the smallest eigenvalues of $X\coloneqq M_{\text{inn}}\hat{A}_{i}$. Hence, by computing a low-rank approximation of $X$:
(32) $X\simeq U_{k}\Lambda_{k}V_{k}^{T},$
we can obtain the following truncated eigendecomposition of $Y$:
(33) $Y\simeq U_{k}(\mathbb{I}_{k}-\Lambda_{k})V_{k}^{T},$
which gives $\Sigma_{k}=\mathbb{I}_{k}-\Lambda_{k}$ and, consequently:
(34) $\Theta_{k}=\mathbb{I}_{k}+\left(\mathbb{I}_{k}-\Lambda_{k}\right)+\left(\mathbb{I}_{k}-\Lambda_{k}\right)\left(\mathbb{I}_{k}-\left(\mathbb{I}_{k}-\Lambda_{k}\right)\right)^{-1}\left(\mathbb{I}_{k}-\Lambda_{k}\right)=\Lambda_{k}^{-1},$
therefore matching the largest eigenvalues of the correction with the smallest eigenvalues of the preconditioned subsystem.
Ultimately, applying the above procedure to each of the $2^{s}$ subsystems results in the following low-rank corrected AMG, henceforth denoted as LRCAMG($k$):
(35) $\mathbb{I}_{2^{s}}\otimes M_{\text{inn}}+\begin{pmatrix}U_{k,1}\Theta_{k,1}V_{k,1}^{T}&&\\ &\ddots&\\ &&U_{k,2^{s}}\Theta_{k,2^{s}}V_{k,2^{s}}^{T}\end{pmatrix}.$
LRCAMG($k$) is compatible with SpMM in the application of $M_{\text{inn}}$. However, this comes at the price of introducing another level of approximation by ignoring $\hat{A}$’s outer-couplings. While low-rank corrections proved very effective on LRCFSAI($k$), even accelerating its convergence, (relatively) low-rank perturbations had a critical impact on AMG. Indeed, Table 2 summarises the results obtained with LRCAMG($k$) on the model problem of Eq. 25. Unfortunately, low-rank corrections cannot restore the effectiveness of AMG, which excels in removing error components on the lower part of the spectrum with high accuracy, something that low-rank correction cannot match. Due to its unsatisfactory performance, we abandoned its parallel implementation and explored the better alternatives of Section 3.
Table 2: GMRES + LRCAMG($k$) results on the model problem with $n=64^{3}$.
iterations
preconditioner $s=0$ $s=1$ $s=2$ $s=3$
AMG 6 6 6 6
LRCAMG(0) 6 23 29 29
LRCAMG(1) 9 23 26 24
LRCAMG(2) 9 23 22 24
LRCAMG(4) 9 19 21 20
LRCAMG(8) 9 17 18 18
LRCAMG(16) 9 15 16 15
3 Inner-interface spatial discretisation
This section explores two new approaches to accelerate AMG by making it compatible with SpMM. Unlike LRCFSAI($k$) and LRCAMG($k$), the preconditioners developed here rely on a different ordering of unknowns, henceforth denoted as inner-interface ordering. To define it, let us recall the single symmetry case of Fig. 2(a) and classify all the grid points into interface: those coupled with other subdomain unknowns, and inner: those that are not. Then, the inner-interface ordering arises from applying the mirrored ordering of Fig. 1 to the inner unknowns first and the interface ones afterwards. Remarkably enough, AMGR, the preconditioner derived in Section 3.2, relaxes this constraint, allowing any ordering of the interface unknowns. Additionally, boundary conditions are no longer required to be symmetric, therefore being compatible with repeated geometries like the illustrative Fig. 2(b).
(a) Mirrored geometry
(b) Repeated geometry
Figure 2: 1D meshes with a random inner-interface ordering.
As a result of using the inner-interface ordering, the discrete Laplacian reads:
(36) $A=\begin{pmatrix}\bar{K}&\bar{B}\\ \bar{B}^{T}&\bar{C}\end{pmatrix}\in\mathbb{R}^{n\times n},$
where $\bar{K}\in\mathbb{R}^{n_{\text{inn}}\times n_{\text{inn}}}$, $\bar{B}\in\mathbb{R}^{n_{\text{inn}}\times n_{\text{ifc}}}$, and $\bar{C}\in\mathbb{R}^{n_{\text{ifc}}\times n_{\text{ifc}}}$. Additionally, thanks to the mirrored ordering imposed on the inner and interface grid points, we have that $\bar{K}=\mathbb{I}_{2^{s}}\otimes K$ and $\bar{B}=\mathbb{I}_{2^{s}}\otimes B$. On the other hand, $C$ satisfies Eq. 6, and consequently, SpMM can be exploited to accelerate all the submatrix products while reducing their setup costs and memory footprint.
3.1 Schur complement-based AMG
Given the poor results obtained with LRCAMG($k$), our aim now is to exploit the structure of $\bar{K}$, $\bar{B}$ and $\bar{C}$ to accelerate the AMG approximation of $A$, $M_{A}\simeq A^{-1}$. By virtue of Eq. 36, the inverse of $A$’s LDU factorisation reads:
(37) $A^{-1}=\begin{pmatrix}\mathbb{I}&-\bar{K}^{-1}\bar{B}\\ &\mathbb{I}\end{pmatrix}\begin{pmatrix}\bar{K}^{-1}&\\ &\bar{S}^{-1}\end{pmatrix}\begin{pmatrix}\mathbb{I}&\\ -\bar{B}^{T}\bar{K}^{-1}&\mathbb{I}\end{pmatrix},$
where $\bar{S}\coloneqq\bar{C}-\bar{B}^{T}\bar{K}^{-1}\bar{B}$ is the Schur complement of $\bar{K}$. Then, we can derive valid preconditioners by seeking approximations of $\bar{K}^{-1}$ and $\bar{S}^{-1}$. In our case, given that $\bar{K}^{-1}=\mathbb{I}_{2^{s}}\otimes K^{-1}$, we will approximate $\bar{K}^{-1}$ through the AMG of $K$, $M_{K}\simeq K^{-1}$. This has two major advantages: on the one hand, the setup costs and memory footprint of $M_{K}$ are $2^{s}$ times smaller than those of $M_{\bar{K}}$. On the other, its application is compatible with SpMM.
Regarding the approximation of $\bar{S}^{-1}$, the better it is, the closer the action of our Schur-based AMG will be to that of $M_{A}$. Given the favourable results of Section 2.1, we will build an approximation of $\bar{S}^{-1}$ based on LRCFSAI($k$). For simplicity, let us assume a single reflection symmetry. As before, applying a mirrored ordering to the interface grid points ensures that $\bar{C}$ satisfies the following structure:
(38) $\bar{C}=\begin{pmatrix}C_{1}&C_{2}\\ C_{2}&C_{1}\end{pmatrix},$
where $C_{1},C_{2}\in\mathbb{R}^{n_{\text{ifc}}/2\times n_{\text{ifc}}/2}$ account for the interface-interface couplings of the base mesh with itself and its mirroring, respectively. On the other hand, we have that:
(39) $\bar{S}\coloneqq\bar{C}-\bar{B}^{T}\bar{K}^{-1}\bar{B}=\bar{C}-\mathbb{I}_{2^{s}}\otimes B^{T}K^{-1}B=\begin{pmatrix}S_{1}&C_{2}\\ C_{2}&S_{1}\end{pmatrix},$
where we defined $S_{1}\coloneqq C_{1}-B^{T}K^{-1}B$. Then, in order to compute $\bar{S}^{-1}$ we can proceed as in Eqs. 9, 10, and 11 to block diagonalise the Schur complement of $\bar{K}$:
(40) $\hat{S}\coloneqq Q_{1}\bar{S}Q_{1}^{-1}=\begin{pmatrix}S_{1}+C_{2}&\\ &S_{1}-C_{2}\end{pmatrix}.$
The fact that $\bar{S}$ contains $K^{-1}$ prevents us from using AMG to approximate $\bar{S}^{-1}$. However, given the excellent results obtained with LRCFSAI($k$), it makes sense to mimic its strategy. In particular, we will first build a rough approximation of $\bar{S}^{-1}$ based on an FSAI of $S_{1}$ and compatible with SpMM. Then, we will correct it by means of low-rank corrections to make it approach the (block diagonal version of the) most accurate Schur complement at hand:
(41) $\bar{S}_{\text{AMG}}\coloneqq\bar{C}-\mathbb{I}_{2^{s}}\otimes B^{T}M_{K}B,$
but whose inverse is unavailable.
With this aim, let $L_{K}$ be the lower Cholesky factor of $K$, and $G_{K}\simeq L_{K}^{-1}$ its lower FSAI factor. Then, we have that $G_{K}^{T}G_{K}\simeq K^{-1}$, and it is meaningful to approximate $S_{1}^{-1}$ through an FSAI of $C_{1}-B^{T}G_{K}^{T}G_{K}B\simeq S_{1}$:
(42) $S_{1}^{-1}\simeq G_{S_{1}}^{T}G_{S_{1}}.$
Finally, by recalling Eq. 40 and the fact that $S_{1}$ is substantially denser than $C_{2}$, we can build the sought approximation of $\hat{S}^{-1}$ as follows:
(43) $\hat{S}^{-1}\simeq\mathbb{I}_{2}\otimes G_{S_{1}}^{T}G_{S_{1}},$
which not only is compatible with SpMM but also allows invoking Theorem 2.1 on each decoupled block:
(44) $\hat{S}_{\text{LRC}}^{-1}\coloneqq\mathbb{I}_{2}\otimes G_{S_{1}}^{T}G_{S_{1}}+\begin{pmatrix}Z_{k}^{\text{(1)}}\Theta_{k}^{\text{(1)}}{Z_{k}^{\text{(1)}}}^{T}&\\ &Z_{k}^{\text{(2)}}\Theta_{k}^{\text{(2)}}{Z_{k}^{\text{(2)}}}^{T}\end{pmatrix}.$
Regardless of how we approximate $\bar{S}^{-1}$, Eq. 37 would require applying $M_{K}$ twice, whereas $M_{A}$ is only applied once. In order to avoid such an expensive overhead, we can define our Schur-based AMG, henceforth denoted AMGS, as follows:
(45) $M_{\text{Schur}}\coloneqq\begin{pmatrix}\mathbb{I}&-G_{\bar{K}}^{T}G_{\bar{K}}\bar{B}\\ &\mathbb{I}\end{pmatrix}\begin{pmatrix}\text{AMG}_{\bar{K}}&\\ &Q_{s}^{-1}\hat{S}_{\text{LRC}}^{-1}Q_{s}\end{pmatrix}\begin{pmatrix}\mathbb{I}&\\ -\bar{B}^{T}M_{\bar{K}}&\mathbb{I}\end{pmatrix},$
where we compacted the notation of $M_{\bar{K}}\coloneqq\mathbb{I}_{2^{s}}\otimes M_{K}$ and replaced its extra application with the FSAI of $\bar{K}$, also compacted to $G_{\bar{K}}^{T}G_{\bar{K}}\coloneqq\mathbb{I}_{2^{s}}\otimes G_{K}^{T}G_{K}$. Remarkably enough, this substitution did not deteriorate the convergence of AMGS significantly, but made it nonsymmetric and, therefore, incompatible with PCG.
The main advantages of AMGS are the fact that its matrix multiplications are compatible with SpMM, together with the smaller memory footprint and setup costs that this allows. Table 3 summarises the results obtained with AMGS on the model problem of Eq. 25. As shown, the combination of FSAI and low-rank corrections was not enough to substantially improve its performance with respect to LRCAMG($k$). Consequently, we discarded its parallel implementation and developed the better alternative of Section 3.2.
Table 3: GMRES + AMGS results on the model problem with $n=64^{3}$.
iterations
preconditioner $s=0$ $s=1$ $s=2$ $s=3$
AMG 6 6 6 6
AMGS 6 21 23 22
3.2 Multigrid reduction
As the model problem of Eq. 25 confirmed, the standard AMG algorithm cannot take advantage of spatial symmetries and needs to be appropriately adapted. In particular, we will develop an AMG reduction framework, hereafter denoted as AMGR, that leverages reflection, translational and rotational symmetries regardless of the boundary conditions. Remarkably enough, AMGR only uses SpMM on the inner unknowns, therefore allowing any ordering of the interface. It relies on a prolongation, $P$, with a block structure similar to Eq. 36:
(46) $P=\begin{pmatrix}\bar{W}\\ \mathbb{I}_{n_{c}}\end{pmatrix}\in\mathbb{R}^{n\times n_{c}},$
where $\bar{W}\in\mathbb{R}^{n_{f}\times n_{c}}$ and $\mathbb{I}_{n_{c}}\in\mathbb{R}^{n_{c}\times n_{c}}$. Additionally, $n_{f}$ and $n_{c}$ correspond to the number of fine and coarse nodes as defined by the multigrid reduction.
Ideally, to have the fastest possible coarsening, we should only identify as coarse the interface unknowns, i.e., the $n_{\text{ifc}}$ unknowns associated with the interface block, $\bar{C}$. However, realistically sized problems make it impossible to accurately interpolate inner unknowns solely using interface values. This is due to the large connection distance that may occur. Indeed, if there are $n_{\text{inn}}$ inner unknowns, the maximum inner-interface distance will be of the order of $d]{n_{\text{inn}}}$, where $d$ stands for the geometrical dimension of the problem. Then, a 3D problem with about a million inner unknowns would result in distances of about 100 units, which exceeds the applicability of long-distance interpolation formulas such as Extended+I (ExtI) [18] or dynamic-pattern Least Squares Fit (LSF) [43].
Hence, to allow for an accurate interpolation, we need to convert some inner nodes into coarse. In principle, we can apply any standard coarsening strategy to the inner part of each subdomain’s sub-block. That is, by choosing a strength of connection measure, e.g., a classical strength measure, filtering the resulting adjacency graph to obtain $T$, and running a maximum independent set (MIS) algorithm on $T$ to select which inner nodes become coarse. However, even if many connections are retained in $T$, the number of coarse nodes generally becomes too high, making the resulting operator’s complexity impractical. To cope with this issue, we use an aggressive coarsening strategy designed ad hoc for this problem. Instead of selecting independent nodes from $T$, we select independent nodes from $T^{k}$, for a small power $k$. Using $T^{k}$ is equivalent to constructing an independent set by considering dependent to each other all nodes at a distance smaller or equal to $k$. Larger distances between coarse nodes are not an issue if long-distance interpolations are used. ExtI interpolation is most effective when $k$ is limited to 2. Conversely, LSF allows for larger values of $k$ and coarser operators.
Table 4 summarises the results obtained with AMGR on the model problem of Eq. 25. Adopting the aggressive coarsening that symmetries and repeated geometries induce does not harm convergence and allows for a small and sparse coarse level operator, $A_{c}=P^{T}AP$. Then, if $A_{c}$ is not small enough, we can approximate its inverse using a standard AMG. Additional advantages of AMGR include the compatibility with SpMM of the FSAI used as $\bar{K}$-smoother, reducing its memory footprint and setup costs.
Table 4: PCG + AMGR results on the model problem with $n=64^{3}$.
iterations
preconditioner $s=0$ $s=1$ $s=2$ $s=3$
AMG 6 6 6 6
AMGR 6 9 9 9
Given the promising results obtained with AMGR, we tackled its parallel implementation, which is discussed in Section 4, and tested it in industrial CFD applications in Section 5.
4 Practical implementation
Discretising complex geometries becomes simpler thanks to exploiting spatial symmetries. Indeed, the strategies presented only require meshing the base mesh, which, assuming $n_{b}$ subdomains, corresponds to a $1/n_{b}$ fraction of the entire domain. Then, the implementation expands the base mesh by imposing a symmetry-aware ordering and leveraging the resulting structure of the operators (see Eq. 6). Hence, it is not necessary to build exactly symmetric meshes, and a significant amount of memory and computational resources are saved.
Similarly, to replace SpMV with SpMM effectively, it is required to apply a consistent domain partitioning. Namely, to distribute the base mesh among the available computing resources and extend such a partitioning to the remaining subdomains by the symmetries. Figure 3 illustrates the above procedure on an arbitrary 2D grid.
Figure 3: Adequate partitioning of a mesh with 2 reflection symmetries.
The proposed methods have been implemented on top of Chronos [30], a sparse linear algebra library designed for parallel computers. Chronos provides iterative solvers for linear systems and eigenproblems and advanced preconditioners based on approximate inverses and AMG. The library is written in C++ with a strongly object-oriented design to ease its development and maintenance. It presents a hybrid programming model using MPI for inter-node communication and OpenMP and CUDA to take advantage of manycore processors and GPU accelerators, respectively.
As to the preconditioners, Chronos provides both static and adaptive pattern FSAI (sFSAI or aFSAI), which can be used as standalone preconditioners or as smoothers within a multigrid hierarchy. Chronos also provides an AMG preconditioner implementing a classical coarsening (that is, a division of the nodes into fine and coarse with no aggregation) and several interpolation schemes. Namely, classical and ExtI for Poisson-like problems and LSF for elasticity problems. Additionally, AMG’s quality is improved by using energy minimisation [32] and its application cost reduced through prolongation filtering.
5 Experimental results
This section investigates the advantages of using AMGR to leverage reflection and translational symmetries in industrial applications. All the CFD cases considered are governed by the incompressible Navier-Stokes and the continuity equations:
(47) $\frac{\partial u}{\partial t}+(u\cdot\nabla)u=\nu\Delta u-\frac{1}{\rho}\nabla p\;\;\;\;\;\;\;\nabla\cdot u=0,$
where $\rho$ and $\nu$ are the density and kinematic viscosity, and $u$ and $p$ are the velocity and pressure fields, respectively.
Regardless of the strategy applied for solving the pressure-velocity coupling of Eq. 47, a Poisson equation arises, and its solution represents the most computationally intensive part of the simulations. Considering a classical fractional step projection method [17], at each time iteration, one computes a predictor velocity, $u^{*}$, to later project it onto a divergence-free space. This is done through the gradient of pressure, which is obtained by solving the following Poisson equation:
(48) $\Delta p=\frac{\rho}{\Delta t}\nabla\cdot u^{*},$
where $\Delta t$ is the time-step following the time integration. See [50, 56] for further details about the discretisation employed.
The numerical experiments consist of solving Eq. 48 using three industrial test cases: the flow around a realistic car model, a finned-tube heat exchanger, and the simulation of a wind farm. While the first two allow for exploiting reflection symmetries, the last allows for studying the further advantages of translational and rotational symmetries. For simplicity, the right-hand side (RHS) used in the tests is a random vector, which ensures a broad spectrum of frequencies in the resulting residual, allowing, in turn, for a complete assessment of the preconditioner’s effectiveness in damping all error components. All the executions rely on combined MPI and multithreaded parallelism and have been conducted on the JFF cluster at the Heat and Mass Transfer Technological Center. Its non-uniform memory access (NUMA) nodes are equipped with two Intel Xeon 6230 CPUs (20 cores, 2.1 GHz, 27.5 MB L3 cache and 140 GB/s memory bandwidth) linked to 288GB of RAM and interconnected through 7 GB/s FDR Infiniband.
Figure 4: Industrial test cases used in the numerical experiments. Top: DrivAer car model and finned-tube heat exchanger. Bottom: $6\times 4$ wind farm.
The coarsening ratio and average nonzeros per row in Tables 5, 6, and 7 correspond to the operator at the first level of the multigrid hierarchy, i.e., after the first coarsening, which, in the case of AMGR, is induced by the reduction framework. Additionally, $n_{b}$ denotes the number of subdomains in which the domain can be decomposed, i.e., the number of repeated blocks in $\bar{K}=\mathbb{I}_{n_{b}}\otimes K$.
At the top of Fig. 4, there are displayed the two test cases considered for assessing the performance of AMGR when exploiting reflection symmetries. Namely, the simulation of the flow around the DrivAer fastback car model [27] and within a finned-tube heat exchanger, on which we exploit one and three reflection symmetries, respectively. The results for the DrivAer case in Table 5 make apparent the advantages of AMGR, which preserves the excellent convergence of the standard AMG and, thanks to the faster top-level smoothing granted by replacing SpMV with SpMM, yields 43% speed-ups. The larger the number of repeated blocks in $\bar{K}=\mathbb{I}_{n_{b}}\otimes K$, the greater the benefits of replacing SpMV with SpMM. In this sense, the heat exchanger problem extends the previous results up to 3 symmetries. According to Table 6, regardless of $n_{b}$, the aggressive coarsening induced by the multigrid reduction results in almost the same coarse operator. Indeed, both the coarsening rate and density remain roughly constant. Nevertheless, larger $n_{b}$ implies faster top-level smoothing, which grants up to 68% accelerations.
Table 5: DrivAer problem with 106.4M unknonws on five JFF nodes.
preconditioner $n_{b}$ coarsening ratio avg nnzr its t-sol (s) speed-up
AMG 1 0.36 14.7 26 7.71 1.00
AMGR 2 0.14 37.4 26 5.39 1.43
Table 6: Heat exchanger problem with 18.4M unknonws on two JFF nodes.
preconditioner $n_{b}$ coarsening ratio avg nnzr its t-sol (s) speed-up
AMG 1 0.36 14.5 20 1.54 1.00
AMGR 2 0.14 37.5 19 1.12 1.38
AMGR 4 0.15 37.4 19 1.03 1.50
AMGR 8 0.15 37.6 18 0.91 1.68
The results in Tables 5 and 6 make clear the potential of applying AMGR on domains arising from recurring structures, such as the $6\times 4$ wind farm in Fig. 4. Indeed, while reflection symmetries are generally restricted to $n_{b}\leq 8$, rotational and translational symmetries are not, allowing, for instance, the simulation of an arbitrarily large wind farm by just discretising a single wind turbine. Table 7 summarises the results obtained on a $6\times 4$ wind farm discretised according to [15]. The wind turbines are introduced through the immersed boundary method and, therefore, discretised with a structured grid stretched around the blades and coarsened vertically towards the atmospheric boundary layer. The first thing to note is that the optimal number of subdomains into which dividing the wind farm is not $n_{b}=24$. In fact, relatively small values of $n_{b}$ led to maximum speed-ups, as the application of AMG on the reduced operator quickly counterbalanced SpMM’s accelerations. This follows from the fact that the larger $n_{b}$, the faster SpMM, but also the larger $n_{\text{ifc}}$. Then, according to Eq. 46, the less effective the multigrid reduction and, as observed in Table 7, the larger the coarsening ratio. The wind farm problem is susceptible to such a trade-off due to the reduction itself being less aggressive, as the larger coarsening ratios indicate. Nevertheless, AMGR still proved effective, being up to 70% faster than the standard AMG.
The coefficient matrices arising from the industrial CFD applications considered are remarkably sparse. As a result, so is the optimal top-level smoother, and the best results were obtained using a very light FSAI. Despite obtaining significant speed-ups, this compromised the advantages of AMGR. Not only by making SpMM’s acceleration quickly counterbalanced by the application of AMG on the reduced operator, but also by making both AMG and AMGR have comparable memory footprints. In this sense, higher-order schemes or, more generally, applications entailing a denser Poisson’s equation (e.g., linear elasticity or geomechanical problems) would strengthen the benefits of SpMM and, therefore, of AMGR.
Table 7: Wind farm problem with 52M unknonws on two JFF nodes.
preconditioner $n_{b}$ coarsening ratio avg nnzr its t-sol (s) speed-up
AMG 1 0.41 13.6 37 11.0 1.00
AMGR 2 0.23 22.1 35 8.06 1.36
AMGR 3 0.23 22.0 32 6.48 1.70
AMGR 4 0.25 20.4 36 7.54 1.46
AMGR 6 0.24 21.7 38 8.50 1.29
AMGR 8 0.25 20.2 38 8.74 1.26
AMGR 12 0.25 20.7 37 8.78 1.25
AMGR 24 0.26 19.9 42 11.0 1.00
When it comes to the scalability of AMGR, it is worth recalling Section 4. As discussed, to replace SpMV with SpMM, it is necessary to apply a consistent domain partitioning. Then, instead of distributing the rows of $\bar{K}=\mathbb{I}_{n_{b}}\otimes K$ among all the available resources, we need to distribute the smaller $K$ so that its coefficients are effectively reused in the SpMM (analogously for the top-level smoother). Of course, distributing the smaller sub-matrix entails potentially larger communications. According to the results of Fig. 5, this is not a severe issue. On the one hand, AMGR preserves the excellent weak scalability of AMG. On the other, even if AMGR’s strong scalability is affected by the extra communication overheads, this effect is far from critical, especially considering the relatively large workloads that extreme-scale simulations typically entail.
Figure 5: AMG and AMGR scaling on the unit cube. Left: weak scaling with a workload of $384^{3}$ unknowns per node. Right: strong scaling on a $256^{3}$ mesh.
6 Conclusions
This paper presented a multigrid reduction framework for accelerating AMG on regular domains. Given that modern supercomputers prioritise FLOP performance, most computational physics applications are memory-bound and, therefore, unable to achieve hardware’s theoretical peak performance. To address this limitation, we first showed that given an arbitrarily complex geometry presenting reflection, translational or rotational symmetries, it is possible to apply a consistent ordering that makes the coefficient matrix (and preconditioners) satisfy a regular block structure. This, in turn, allows the standard SpMV to be replaced with the more compute-intensive SpMM, accelerating the application of the preconditioners. This paper focused on accelerating AMG owing to its numerical and computational effectiveness. However, it is worth noting that the intermediate strategies presented naturally apply to other preconditioners.
The proposed multigrid reduction framework introduces an aggressive coarsening to the multigrid hierarchy, reducing the memory footprint, setup and application costs of the top-level smoother. While preserving the excellent convergence of AMG, replacing SpMV with SpMM yielded significant speed-ups. Remarkably enough, AMGR, the resulting preconditioner, does not have any specific requirements on the boundary conditions, only on the mesh and its ordering. Numerical experiments on industrial CFD applications demonstrated up to a 70% speed-up in the solution of Poisson’s equation by AMGR compared to AMG. Furthermore, even if AMGR entail higher communication overheads, strong and weak scalability analyses revealed no serious degradation compared to AMG.
The coefficient matrices arising from the incompressible CFD applications considered are remarkably sparse and, as a result, so is the optimal top-level smoother. Indeed, the best results were obtained smoothing with a very light FSAI and, even if we obtained significant speed-ups, this compromised the advantages of AMGR. Firstly, leveraging relatively few symmetries led to maximum speed-ups, as SpMM’s accelerations were quickly counterbalanced by the application of AMG on the reduced operator, whose size grows slightly with the number of subdomains. Secondly, both AMG and AMGR had a comparable memory footprint. Again, this follows from the particularly lightweight smoother required by the cases considered. In this sense, higher-order schemes or, more generally, applications entailing a denser Poisson’s equation (e.g., linear elasticity or geomechanical problems) would strengthen the advantages of SpMM and, therefore, AMGR.
For all these, immediate lines of work include applying AMGR to (denser) problems arising from structural mechanics. Additionally, we plan to optimise the setup phase, which was irrelevant for the incompressible CFD applications considered but is crucial for simulations with variable Poisson’s equation. Significant accelerations are expected from computing the top-level smoother on the base mesh instead of the entire domain. Finally, we want to tackle the GPU implementation of AMGR and study the extension of the multigrid reduction framework towards nonregular domains, which, despite not exploiting SpMM, would benefit from the aggressive coarsening.